### v1

# Detecting Collective Excitations in Self-Gravitating Bose-Einstein Condensates via Faraday Waves

Abstract
We propose Faraday waves as a probe for collective excitations in self-gravitating Bose-Einstein condensates (SGBECs), driven by periodic modulation of the $s$-wave scattering length. Linear stability analysis of the driven Gross-Pitaevskii-Newton equations reveals that parametric instability follows a Mathieu-like equation, with Faraday waves emerging resonantly when half the driving frequency matches the collective excitation frequency of the SGBEC. This framework yields a stability phase diagram demonstrating the coexistence and interplay between the intrinsic Jeans instability of self-gravitating systems and the unstable tongues of parametric resonance. The phase diagram shows that reducing the driving frequency fragments and narrows parametric resonance structures while expanding the Jeans-dominated regime. We further identify a thermal-gravity competition in the Faraday wavevector, manifested as a sharp transition when thermal energy balances mean-field interaction at low driving frequencies. We numerically simulate Faraday wave formation and dynamics within the SGBEC, including the effects of dissipation; simulations reveal that Faraday wave patterns exhibit high sensitivity to Jeans frequency, transitioning from parametric resonance to Jeans collapse as gravitational strength increases.
I Introduction
Gravity, historically the first fundamental interaction to be characterized, presents unresolved challenges in quantum regimes. Persistent uncertainty surrounds its fundamental nature: whether gravity is quantum or classical Oppenheim (2023); Kryhin and Sudhir (2025). This uncertainty persists because the extreme weakness of gravitational interactions prevents definitive experimental tests of quantum gravitational behavior. Consequently, analog quantum systems that simulate gravitational phenomena under controlled conditions offer a vital pathway forward. These systems engineer long-range attractive interactions that simulate key aspects of gravity. This approach provides crucial insights into collective quantum behavior under gravitational-like potentials. It also helps distinguish intrinsic gravitational features from those of other fundamental interactions Barcelo et al. (2011); Liang et al. (2024a); Marletto and Vedral (2025).
Within this context, self-gravitating quantum many-body systems present a particularly intriguing but challenging subject Chavanis (2002); Andersson and Schmidt (2009); Großardt (2016); Ourabah (2020, 2023); Asakawa et al. (2024). Yet even foundational aspects, such as their collective excitation spectra, remain topics of debate Chavanis (2017); Akbari-Moghanjoughi (2024). Recent advances in ultracold atomic physics, however, offer an excellent solution. The exceptional purity and unprecedented tunability of ultracold atomic gases make them an ideal platform for quantum simulation Morsch and Oberthaler (2006); Bloch et al. (2008). Crucially, the collective behavior of self-gravitating systems, akin to ensembles with long-range attractive interactions, can be engineered in these gases. This can be experimentally achieved by inducing effective gravitational attraction through the application of non-resonant, intense optical dipole forces from laser beams O’dell et al. (2000); Giovanazzi et al. (2001); Choi (2002), or through two contra-propagating quasiresonant lasers Chalony et al. (2013).
Building on these experimental capabilities, we focus specifically on self-gravitating Bose-Einstein condensates (SGBECs). SGBECs are governed by the Gross-Pitaevskii-Newton (GPN) equations, which couple the nonlinear Schrödinger equation describing the condensate wavefunction to the Poisson equation for the self-consistent gravitational potential Paredes et al. (2020). The GPN framework has enabled extensive analysis of SGBECs, including nonlinear dynamics Cartarius et al. (2008), equilibrium configurations Guzman and Urena-Lopez (2006), and quantized vortex formation Asakawa and Tsubota (2024). Beyond their role as laboratory analogs, SGBECs are of significant interest in astrophysics and cosmology Hu et al. (2000); Boehmer and Harko (2007); Jones and Bernstein (2001); Schroven et al. (2015); Chavanis (2020); Guo et al. (2021); Kirejczyk et al. (2022); Sivakumar et al. (2025). They serve as models for ultralight bosonic dark matter candidates Chavanis (2011a, b, 2014); Suárez et al. (2014); Chavanis (2021); Nikolaieva et al. (2021, 2023), potentially explaining core structures in galactic halos, and for boson stars Kolb and Tkachev (1993); Chavanis and Harko (2012); Dmitriev et al. (2021); Levkov et al. (2017); Dmitriev et al. (2024), hypothetical compact objects composed of scalar fields. Understanding the dynamics and collective excitations of SGBECs is thus crucial for probing fundamental physics across scales.
To probe the collective excitations of SGBECs, the generation and observation of Faraday waves provide a powerful diagnostic tool Hernández-Rajkov et al. (2021); Cominotti et al. (2022). Faraday waves are striking standing wave patterns that emerge on a fluid surface subjected to parametric driving, i.e., when an external control parameter (like vertical oscillation amplitude) is varied periodically. First documented by Michael Faraday in 1831 Faraday (1831), this phenomenon represents a spontaneous breaking of both temporal and spatial symmetry. Significantly, Faraday waves have been investigated in diverse quantum systems Staliunas et al. (2002, 2004); Nicolin et al. (2007); Bhattacherjee (2008); Nath and Santos (2010); Nicolin and Raportaru (2010); Pollack et al. (2010); Nicolin (2011); Capuzzi et al. (2011); Balaž and Nicolin (2012); Łakomy et al. (2012); Raportaru (2012); Balaž et al. (2014); Abdullaev et al. (2015); Sudharsan et al. (2016); Nguyen et al. (2019); Vudragović and Balaž (2019); Abdullaev et al. (2019, 2019); Turmanov et al. (2020); Liu and Tu (2020); Cheng and Shi (2021); Zhang et al. (2020); Maity et al. (2020); Kwon et al. (2021); Zhang et al. (2022); Shukuno et al. (2023); Brito et al. (2023); Liang et al. (2024b); Wan et al. (2024); del Río-Lima et al. (2024); Wang et al. (2025), including their first experimental demonstration in atomic BECs by Engels et al. Engels et al. (2007) and subsequent studies in dipolar condensates Łakomy et al. (2012); Turmanov et al. (2020), two-component BECs Chen et al. (2019); Wang et al. (2025), spin-orbit-coupled BECs Zhang et al. (2022) and Rydberg-dressed BECs Shi et al. (2023). These theoretical analyses and experiments establish that Faraday waves directly probe the underlying collective excitation spectrum of the quantum fluid. Therefore, driving Faraday waves in SGBECs offers a novel and controlled framework for investigating the collective dynamics, particularly the excitation spectrum of quantum fluids dominated by long-range attractive interactions.
In this paper, we present a theoretical investigation of Faraday waves in SGBECs, driven by periodic modulation of the $s$-wave scattering length. We employ linear stability analysis of the GPN equations to derive a Mathieu-like equation governing the parametric instability. This allows us to map out the stability phase diagram, revealing the interplay between the unstable tongues characteristic of parametric resonance and the Jeans instability inherent to self-gravitating systems. We analyze the dependence of the Faraday wavevector on the driving parameters and self-gravity. Furthermore, we incorporate dissipation effects and present numerical simulations illustrating the formation and evolution of Faraday waves within the SGBEC.
The paper is structured as follows: Section II outlines the theoretical model, the GPN equation with time-dependent interaction strength, and details the linear stability analysis leading to the Mathieu-like equation. Section III presents the resulting stability phase diagrams and analyzes the characteristics of the Faraday wavevector. Section IV discusses numerical simulations of Faraday wave dynamics within the driven SGBEC, including the effects of dissipation. We conclude in Section V.
II Model and Linear stability analysis
II.1 The Gross-Pitaevskii-Newton equation
The dynamics of self-gravitating Bose-Einstein condensates (SGBECs) is governed by the coupled Gross-Pitaevskii-Newton(Possion) (GPN or GPP) equations:
${\rm i}\hbar\frac{\partial\Psi}{\partial t}$ $=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\rm ext}\Psi+g(t)|\Psi|^{2}\Psi+m\Phi\Psi,$ (1)
$\nabla^{2}\Phi$ $=4\pi Gm|\Psi|^{2},$
where $\Psi(\mathbf{r},t)$ is the condensate wave function, $V_{\rm ext}$ denotes the external potential, $m$ is the particle mass, and $\Phi(\mathbf{r},t)$ is the gravitational potential which satisfies the Poisson equation. The nonlinear coupling strength $g(t)=4\pi\hbar^{2}a_{s}(t)/m$ depends on the $s$-wave scattering length $a_{s}(t)$, tunable via Feshbach resonance. The GPN system can be expressed equivalently in integral form:
[…]
Without gravity ($\Phi=0$), Eq. (1) reduces to the standard Gross-Pitaevskii equation; neglecting the mean-field term ($g=0$) yields the Schrödinger-Newton equation.
A generalized GPN model incorporating logarithmic nonlinearity and dissipation is proposed in Chavanis (2017):
${\rm i}\hbar\frac{\partial\Psi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\rm ext}\Psi+g(t)|\Psi|^{2}\Psi+m\Phi\Psi-\beta^{-1}\ln|\Psi|\Psi-{\rm i}\hbar\gamma\Psi,$ (3)
where $\gamma$ is the damping coefficient. The term $-\beta^{-1}\ln|\Psi|^{2}\Psi$ accounts for finite-temperature effects, where $\beta^{-1}=-k_{B}T$ ($k_{B}$ is Boltzmann’s constant, $T$ is temperature), and the last term describes phenomenological dissipation. Eq. (1) is recovered when $\beta^{-1}=0$ and $\gamma=0$.
II.2 Linear stability analysis
Faraday waves in BECs are excited by periodic modulation of either the external potential or the coupling strength $g(t)$ Engels et al. (2007). Here we modulate $g(t)$ as
$g(t)=g_{0}\left(1+\alpha\cos\omega t\right),$ (4)
where $g_{0}=4\pi\hbar^{2}a_{s}/m$ is the unmodulated coupling strength, $\omega$ is the driving frequency, and $\alpha$ is the modulation depth. We consider a cigar-shaped condensate tightly confined in the transverse ($xy$) plane, with driving along the longitudinal ($z$) direction. When the Faraday wavelength $k^{-1}$ is much smaller than the condensate radius $R_{\rm TF}$, the external potential $V_{\rm ext}$ is negligible.
To analyze stability, we decompose $\Psi=\Psi_{B}\Psi_{P}$, where $\Psi_{B}(\mathbf{r}_{\perp})$ is the transverse ground state and $\Psi_{P}(z,t)$ is the longitudinal perturbation. The equilibrium state satisfies Hernández-Rajkov et al. (2021); Hai et al. (2008)
$g_{0}|\Psi_{B}|^{2}+m\Phi_{B}-\beta^{-1}\ln|\Psi_{B}|^{2}=E_{B},$ (5)
with $\nabla^{2}\Phi_{B}=4\pi mG|\Psi_{B}|^{2}$, where $E_{B}$ is the equilibrium energy eigenvalue. The unperturbed dynamics follows
${\rm i}\hbar\frac{\partial\Psi_{B}}{\partial t}=\left[-\frac{\hbar^{2}}{2m}\nabla^{2}_{\perp}+E_{B}\right]\Psi_{B}.$ (6)
Substituting $\Psi=\Psi_{B}\Psi_{P}$ into Eq. (3) and using Eq. (6) yields the perturbation equation:
${\rm i}\hbar\frac{\partial\Psi_{P}}{\partial t}=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}\Psi_{P}}{\partial z^{2}}+g(t)|\Psi_{B}|^{2}\left(|\Psi_{P}|^{2}-1\right)\Psi_{P}+m(\Phi-\Phi_{B})\Psi_{P}-\beta^{-1}\ln|\Psi_{P}|\Psi_{P}-{\rm i}\hbar\gamma\Psi_{P}.$ (7)
Linearizing around $\Psi_{P}=1$ for small perturbations $\Psi_{P}(z,t)=1+W(t)\cos(kz)$ with $W(t)=U(t)+{\rm i}V(t)$ gives coupled equations:
$\hbar\dot{V}$ $=-\epsilon_{k}U-2g(t)|\Psi_{B}|^{2}U+\frac{2m\Omega_{J}^{2}}{k^{2}}U+\beta^{-1}U+\hbar\gamma V,$ (8)
$\hbar\dot{U}$ $=\epsilon_{k}V+\hbar\gamma U,$
where $\epsilon_{k}=\hbar^{2}k^{2}/(2m)$ is the free-particle energy, and $\Omega_{J}=\sqrt{4\pi Gm|\Psi_{B}|^{2}}$ is the Jeans frequency. Combining these yields a damped Mathieu equation:
$\frac{{\rm d}^{2}U}{{\rm d}t^{2}}+2\gamma\frac{{\rm d}U}{{\rm d}t}+\left[\Omega_{k}^{2}+\frac{\alpha g_{0}|\Psi_{B}|^{2}}{m}k^{2}\cos\omega t\right]U=0,$ (9)
with the characteristic excitation frequency
$\Omega_{k}^{2}=\Omega_{\rm Bog}^{2}-\Omega_{J}^{2}.$ (10)
Here $\Omega_{\rm Bog}$ is the Bogoliubov spectrum:
$\Omega_{\rm Bog}^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+c_{s}^{2}k^{2},\quad c_{s}^{2}=\frac{g_{0}n}{m}-\frac{\beta^{-1}}{m},$ (11)
where $n=|\Psi_{B}|^{2}$ is the equilibrium density, and $c_{s}$ is the sound speed. The term $g_{0}n/m$ arises from mean-field interactions, and $-\beta^{-1}/m$ from finite-temperature effects. In the zero-temperature limit ($\beta^{-1}=0$), $c_{s}^{2}=g_{0}n/m$.
Fig. 1 shows the collective excitation spectrum of SGBECs. For large $k$ ($k\xi\gg 1$), where $\xi=\hbar/\sqrt{2mg_{0}n}$ is the healing length, $\Omega_{k}$ approaches the Bogoliubov dispersion. Near $k_{J}\xi\equiv\sqrt{\sqrt{1+\Omega^{2}_{J}/\Omega^{2}_{0}}-1}$ (with $\Omega_{0}=g_{0}n/\hbar$), the Jeans instability causes $\Omega_{k}^{2}<0$, indicating exponential growth. Here $k_{J}$ is the critical wavenumber defining the quantum Jeans scale Chavanis (2020).
Figure 1: (Color online) Collective excitation spectrum of SGBECs. Parameters: $\Omega_{0}=g_{0}n/\hbar$, $\xi=\hbar/\sqrt{2mg_{0}n}$. The region $k<k_{J}$ indicates Jeans instability.
Using the dimensionless time $\tau=\omega t/2$, Eq. (9) becomes
$\frac{d^{2}U}{d\tau^{2}}+\bar{\gamma}\frac{dU}{d\tau}+\left[\lambda-2q\cos(2\tau)\right]U=0,$ (12)
where $\bar{\gamma}=4\gamma/\omega$, and
$\lambda=\frac{4\Omega_{k}^{2}}{\omega^{2}},\quad q=-\frac{2\alpha g_{0}n}{m}\frac{k^{2}}{\omega^{2}}.$ (13)
The parameter $\lambda$ encodes the excitation spectrum of the self-gravitating quantum fluid. For $\gamma=0$, Eq. (12) reduces to the canonical Mathieu equation, with well-established stability properties McLachlan (1947); Abramowitz and Stegun (1965); Wang and Guo (1989). We analyze its phase diagram and derive the Faraday wavevector in the following section.
III Stability and parametric resonance
III.1 Stability phase diagram
According to Floquet theory, the solution of Eq. (12) in the dissipationless limit can be expressed as:
$U(\tau)=e^{\mu\tau}f(\tau),$ (14)
where $f(\tau)$ is a periodic function with period $\pi$. The Floquet exponent $\mu$ determines the stability of solutions. When the real part of $\mu$ equals zero, the solution becomes periodic and corresponds to the boundary between stable and unstable regions. Solutions are unstable when Re$(\mu)>0$ and stable when Re$(\mu)<0$.
Figure 2: (Color online) Stability phase diagram showing primary ($\nu=1$), secondary ($\nu=2$), and third ($\nu=3$) instability tongues. Shaded regions denote unstable solutions.
The numerically computed stability phase diagram is presented in Fig. 2. For small values of $q$, unstable regions emerge near $\lambda=\nu^{2}$ ($\nu=1,2,3,\ldots$), consistent with the characteristic instability tongues of parametric resonance. The primary instability tongue ($\nu=1$) exhibits the largest spatial extent in parameter space, with higher-order tongues ($\nu=2,3,\ldots$) progressively narrowing. For the dominant $\nu=1$ tongue ($\lambda=1$), the resonance condition corresponds to:
$\Omega_{k}=\omega/2,$ (15)
which yields the frequency-matching relation:
$\left(\frac{\omega}{2}\right)^{2}+\Omega_{J}^{2}=\Omega_{\rm Bog}^{2}.$ (16)
Eq.(16) indicates that the Faraday wave only occurs in the SGBEC when the relation Eq.(16) is met.
Fig. 3 illustrates the stability phase diagrams at different Jeans frequencies. Without self-gravity ($\Omega_{J}=0$, Fig. 3(a)), instability arises exclusively from parametric driving. As $\Omega_{J}$ increases (Figs. 3(b)-(d)), the unstable region near $\lambda\approx 0$ expands significantly due to the relationship $\lambda=4\Omega_{k}^{2}/\omega^{2}$. When $\Omega_{\rm Bog}^{2}<\Omega_{J}^{2}$, $\lambda$ becomes negative and its magnitude scales as $|\lambda|\propto\Omega_{J}^{2}$, explaining the expansion of the Jeans-dominated region. Concurrently, the parametric resonance tongues shift toward larger $\lambda$ values. This shift occurs because $\lambda\propto k^{2}$ at intermediate wavevectors (where $\Omega_{k}^{2}\propto k^{2}$), and increased gravitational attraction requires shorter wavelengths to satisfy the parametric resonance condition $\lambda=\nu^{2}$, with the shift magnitude given by $\Delta\lambda\approx(4/\omega^{2})\Delta(\Omega_{J}^{2})$.
Figure 3: (Color online) Stability phase diagrams under different Jeans frequencies with $\Omega_{0}/\omega=0.5$.
Fig. 4 demonstrates the interplay of Jeans instability and parametric resonance at varying driving frequencies $\omega$. At low driving frequencies ($\omega\ll\Omega_{0}$), the instability is dominated by Jeans mechanisms, manifested as a broad unstable band near $\lambda\approx 0$. As $\omega$ decreases, the parametric resonance tongues fragment and narrow significantly due to the scaling $\Delta\lambda\sim\omega^{-2}$ and $\Delta q\sim\omega^{-4}$. Ultimately, the parametric resonance features vanish completely, leaving only the expanding Jeans-dominated region. Conversely, at high $\omega$, distinct parametric resonance tongues stabilize while the Jeans region contracts. The transition between these regimes occurs near $\omega\sim\Omega_{0}$, where competition between the two mechanisms causes the observed tongue fragmentation.
Figure 4: (Color online) Evolution of stability diagrams with decreasing driving frequency $\omega$ at $\Omega_{J}/\Omega_{0}=1$, showing progressive tongue fragmentation and Jeans region expansion.
III.2 The Faraday wavevector
The Faraday wavevector $k_{F}$ is derived from the instability condition $\lambda=\nu^{2}$ :
$k_{F}\xi=\sqrt{\sqrt{\frac{1}{\Omega_{0}^{2}}\left(\Omega_{J}^{2}+\frac{\nu^{2}\omega^{2}}{4}\right)+\left(\frac{\beta^{-1}}{\hbar\Omega_{0}}-1\right)^{2}}+\left(\frac{\beta^{-1}}{\hbar\Omega_{0}}-1\right)}.$ (17)
In the absence of self-gravity and thermal effects ($\Omega_{J}=0$, $\beta^{-1}=0$), this expression reduces to the result for conventional BECs Hernández-Rajkov et al. (2021).
Fig. 5 shows $k_{F}$ as a function of driving frequency $\omega$ for the first three branches. The Faraday wavevector increases with driving frequency, approaching the behavior of non-self-gravitating BECs at large $\omega$. At low driving frequencies, $k_{F}\xi$ approaches a finite value $\sqrt{\sqrt{\Omega_{J}^{2}/\Omega_{0}^{2}+1}-1}$ in SGBECs, contrasting with the $k_{F}\rightarrow 0$ behavior in non-gravitational systems. Higher $\nu$ values yield larger $k_{F}$ at fixed $\omega$, consistent with excitation of shorter-wavelength modes.
Figure 5: (Color online) Faraday wavevector versus driving frequency for $\nu=1,2,3$ with $\Omega_{J}/\Omega_{0}=1.2$ and $\beta^{-1}=0$.
Fig. 6 presents the complete phase diagram of the primary Faraday wavevector ($\nu=1$) as a function of Jeans frequency and $\beta^{-1}$. We observe a sharp increase in $k_{F}$ near $\beta^{-1}/(\hbar\Omega_{0})=1$ at low driving frequencies ($\omega\ll\Omega_{0}$), while the transition remains smooth at high frequencies ($\omega\gg\Omega_{0}$). This arises from the competition between gravitational confinement and thermal fluctuations. At $\delta\equiv\beta^{-1}/(\hbar\Omega_{0})=1$, the thermal energy $k_{B}T$ balances the mean-field interaction energy $g_{0}n$, causing a critical shift in the excitation spectrum. In the low-$\omega$ regime, gravitational dominance amplifies the square-root singularity in Eq. (17), resulting in a steep but continuous rise in $k_{F}$ for $\delta>1$. This signifies a crossover: For $\delta<1$, the system supports stable collective excitations of the SGBEC (Bogoliubov-like modes modified by self-gravity); For $\delta>1$, thermal fluctuations destabilize the condensate, triggering dissipative collapse. High-frequency driving suppresses the singularity via the $\omega^{2}$ term in Eq. (17), yielding a smooth crossover.
Figure 6: (Color online) Phase diagram of $k_{F}\xi$ versus $\beta^{-1}/(\hbar\Omega_{0})$ and $\Omega_{J}/\Omega_{0}$ for $\nu=1$.
IV Numerical results
Our numerical simulations are based on solving the dimensionless damped Mathieu equation governing the perturbation amplitude $U(\tau)$ with fixed parameters $\lambda=1$ and $q=-0.1$, where $\bar{\gamma}\equiv 4\gamma/\omega$ is the dimensionless damping coefficient. This configuration satisfies the primary parametric resonance condition $2\Omega_{k}=\omega$ for $\nu=1$. Fig. 7 presents the dynamical evolution across four dissipation strengths, with each column corresponding to: (a) $\bar{\gamma}=0$ exhibiting exponential growth; (b) $\bar{\gamma}=0.01$ showing persistent instability; (c) $\bar{\gamma}=0.02$ demonstrating suppressed growth; and (d) $\bar{\gamma}=0.05$ displaying overdamped decay. The top row tracks $U(\tau)$, while the middle row shows background-subtracted density $|\Psi_{P}|^{2}-1$ to isolate Faraday waves from the uniform ground state, and the bottom row displays momentum-space density.
Figure 7: (Color online) Density wave dynamics under dissipation strengths $\bar{\gamma}=0$ (a), $0.01$ (b), $0.02$ (c), $0.05$ (d) with $\lambda=1$, $q=-0.1$.
For $\bar{\gamma}=0$ (Fig. 7(a)), parametric instability drives exponential growth of $U(\tau)$ with Floquet exponent $\mu\approx 0.2$, generating pronounced density waves of wavelength $\lambda_{F}=2\pi/k$ and sharp momentum-space peaks at ±$\pm k$. At $\bar{\gamma}=0.01$ (Fig. 7(b)), dissipation fails to suppress the instability. $U(\tau)$ maintains strong exponential growth with minimal amplitude reduction. This confirms that weak dissipation cannot overcome parametric resonance. At $\bar{\gamma}=0.02$ (Fig. 7(c)), the competition between dissipation and parametric driving becomes evident: while unstable evolution persists, amplitude growth is significantly suppressed, resulting in weak density modulations. Under strong dissipation ($\bar{\gamma}=0.05$, Fig. 7(d)), dissipation dominates, causing rapid decay of $U(\tau)$ within $\tau<20$ and complete suppression of density waves. This sequence illustrates the transition from parametric-resonance-dominated ($\bar{\gamma}<0.02$) to dissipation-dominated ($\bar{\gamma}>0.02$) regimes. Gravitational effects influence this competition through the $\lambda$-parameter in Eq. (12), which encodes the $\Omega_{J}$-modified excitation spectrum.
In Fig. 8, we present the evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies $\Omega_{J}/\Omega_{0}=0,0.1,0.5,1.2$, with fixed parameters $\omega/\Omega_{0}=1$, $q=-0.1$, and $\bar{\gamma}=0.01$. The corresponding $\lambda$ values, determined by the relation $\lambda=1-4(\Omega_{J}/\Omega_{0})^{2}$, are $\lambda=1.00$, $0.96$, $0.00$, and $-4.76$ respectively. For no self-gravity ($\Omega_{J}/\Omega_{0}=0$, $\lambda=1$), the system resides at the center of the primary parametric resonance tongue, resulting in exponential growth of $U(\tau)$ despite weak damping ($\bar{\gamma}=0.01$), characteristic of parametric instability that seeds Faraday waves. As $\Omega_{J}/\Omega_{0}$ increases to $0.1$ ($\lambda=0.96$), the system remains within the parametric resonance tongue but approaches its boundary, exhibiting sustained growth of $U(\tau)$ at a reduced rate compared to the $\Omega_{J}=0$ case, with Faraday waves forming at diminished amplitudes due to competition between parametric driving and incipient gravitational effects. At the critical transition point ($\Omega_{J}/\Omega_{0}=0.5$, $\lambda=0$) where $\Omega_{k}=0$, parametric resonance and Jeans instability coexist, leading to irregular, non-exponential growth of $U(\tau)$ that deviates from characteristic Faraday patterns. Finally, in the Jeans-dominated regime ($\Omega_{J}/\Omega_{0}=1.2$, $\lambda=-4.76$), self-gravity overwhelms parametric driving ($\Omega_{k}^{2}<0$), causing rapid exponential growth of $U(\tau)$ characteristic of gravitational collapse that suppresses Faraday wave formation. This progression confirms the phase diagram prediction (Fig. 4) where increasing Jeans instability fragments parametric resonance structures, with the transition at $\Omega_{J}/\Omega_{0}\sim 0.5$ ($\lambda=0$) highlighting the sensitivity of driven quantum fluids to long-range attractive interactions.
Figure 8: Evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies: (a) $\Omega_{J}/\Omega_{0}=0$ ($\lambda=1$), (b) $\Omega_{J}/\Omega_{0}=0.1$ ($\lambda=0.96$), (c) $\Omega_{J}/\Omega_{0}=0.5$ ($\lambda=0$), (d) $\Omega_{J}/\Omega_{0}=1.2$ ($\lambda=-4.76$). Fixed parameters: $\omega/\Omega_{0}=1$, $q=-0.1$, $\bar{\gamma}=0.01$. Exponential growth in (a) signifies parametric instability, while rapid growth in (d) reflects Jeans collapse.
V Conclusion
In this work, we have demonstrated that parametric resonance induced by periodic modulation of the $s$-wave scattering length can generate Faraday waves in self-gravitating Bose-Einstein condensates (SGBECs). Through linear stability analysis of the driven Gross-Pitaevskii-Newton equations, we derived a damped Mathieu equation governing the perturbation dynamics. Floquet analysis revealed a unified stability phase diagram that captures the competitive interplay between parametric resonance instabilities and the intrinsic Jeans instability of self-gravitating systems. The Faraday wavevector was systematically analyzed, with its dependence on gravitational strength and thermal effects characterized through comprehensive phase diagrams. Numerical simulations further illustrated the formation and evolution of density wave patterns under various dissipation conditions and Jeans frequencies. The study reveals that Faraday wave patterns exhibit high sensitivity to gravitational strength, transitioning from parametric-resonance-dominated patterns to Jeans-collapse-dominated behavior as gravitational strength increases.
Notably, while this study employed modulation of the scattering length, Faraday waves may also be excited through periodic modulation of the external potential in similar configurations. Given the significance of the generalized GPN equations in astrophysical contexts, our approach provides a flexible platform for exploring gravitational analogs. Specifically, such systems could model scenarios where external potential fields originate from compact objects like black holes, and coupling constants emerge from self-gravitating dark matter configurations. This framework offers promising pathways for laboratory investigations into gravitational phenomena using quantum degenerate gases.
Acknowledgments
This work was supported by the Open Fund of Key Laboratory of Multiscale Spin Physics (Ministry of Education), Beijing Normal University (Grant No.SPIN2024N03), and the Scientific Research Startup Foundation for High-Level Talents at Anqing Normal University (Grant No.241042). We thank Professor Zhanchun Tu for valuable discussions and insightful comments.