### v2

# Detecting Collective Excitations in Self-Gravitating Bose-Einstein Condensates via Faraday Waves

Abstract
We propose Faraday waves as a probe for collective excitations in self-gravitating Bose-Einstein condensates (SGBECs), driven by periodic modulation of the $s$-wave scattering length. Linear stability analysis of the driven Gross-Pitaevskii-Newton equations reveals that parametric instability follows a Mathieu-like equation, with Faraday waves emerging resonantly when half the driving frequency matches the collective excitation frequency of the SGBEC. This framework yields a stability phase diagram that maps the competitive interplay between the unstable tongues of parametric resonance and the intrinsic Jeans instability. The diagram reveals that increasing the driving frequency compresses the Jeans-unstable region and allows well-separated parametric resonance tongues to dominate, thereby creating a clear regime for observing Faraday waves. Conversely, lowering the driving frequency expands the domain of Jeans instability, which fragments and overwhelms the parametric resonance structures. We numerically simulate Faraday wave formation and dynamics within the SGBEC, including the effects of dissipation; simulations reveal a characteristic transition from parametric-resonance-driven Faraday waves to gravity-dominated Jeans collapse as the Jeans frequency increases.
I Introduction
Gravity, the first fundamental interaction to be characterized, remains the least understood in quantum regimes Oppenheim (2023); Kryhin and Sudhir (2025). The extreme weakness of gravitational interactions prevents direct experimental probes of quantum gravitational phenomena in tabletop experiments. Analog quantum systems that emulate gravitational phenomena through engineered long-range attractive interactions provide a valuable alternative for studying quantum many-body behavior under gravitational-like potentials. These systems offer insights into collective quantum dynamics that may parallel those in real gravitational systems, serving as testbeds for theoretical ideas about quantum behavior in gravitational settings Barcelo et al. (2011); Liang et al. (2024a); Marletto and Vedral (2025).
Within this context, self-gravitating quantum many-body systems have attracted growing interest, particularly as analog platforms for exploring gravitational phenomena in controllable quantum settings Chavanis (2002); Andersson and Schmidt (2009); Großardt (2016); Ourabah (2020, 2023); Asakawa et al. (2024). These systems are of fundamental importance for testing quantum-gravitational phenomena and have wide-ranging implications for modeling astrophysical scenarios. Notably, even foundational issues such as the form of the collective excitation spectrum remain contentious. For instance, a hydrodynamic approach yields a classical-like spectrum Chavanis (2017), while a direct perturbative treatment of the quantum equations predicts ultraviolet divergence Akbari-Moghanjoughi (2024). Resolving these theoretical discrepancies requires experimental access to the excitation spectrum of such systems. Ultracold atomic gases, with their exceptional purity and tunability, offer an ideal platform for quantum simulation of these phenomena Morsch and Oberthaler (2006); Bloch et al. (2008). Through techniques such as optical dipole forces or laser-induced interactions, which can generate long-range attractive interactions analogous to gravity O’dell et al. (2000); Giovanazzi et al. (2001); Choi (2002); Chalony et al. (2013), they enable the emulation of self-gravitating quantum fluids and the study of collective behavior under controlled, gravitational-like potentials.
Building on these experimental capabilities, we focus specifically on self-gravitating Bose-Einstein condensates (SGBECs). SGBECs are governed by the Gross-Pitaevskii-Newton (GPN) equations, which couple the nonlinear Schrödinger equation describing the condensate wavefunction to the Poisson equation for the self-consistent gravitational potential Paredes et al. (2020). The GPN framework has enabled extensive analysis of SGBECs, including nonlinear dynamics Cartarius et al. (2008), equilibrium configurations Guzman and Urena-Lopez (2006), and quantized vortex formation Asakawa and Tsubota (2024). Beyond their role as laboratory analogs, SGBECs are of significant interest in astrophysics and cosmology Hu et al. (2000); Boehmer and Harko (2007); Jones and Bernstein (2001); Schroven et al. (2015); Chavanis (2020); Guo et al. (2021); Kirejczyk et al. (2022); Shukla et al. (2024); Sivakumar et al. (2025). They serve as models for ultralight bosonic dark matter candidates Chavanis (2011a, b, 2014); Suárez et al. (2014); Verma et al. (2021); Chavanis (2021); Nikolaieva et al. (2021, 2023), potentially explaining core structures in galactic halos, and for boson stars Kolb and Tkachev (1993); Chavanis and Harko (2012); Dmitriev et al. (2021); Verma et al. (2022); Levkov et al. (2017); Dmitriev et al. (2024), hypothetical compact objects composed of scalar fields. Understanding the dynamics and collective excitations of SGBECs is thus crucial for probing fundamental physics across scales.
To probe the collective excitations of SGBECs, the generation and observation of Faraday waves provide a powerful diagnostic tool Hernández-Rajkov et al. (2021); Cominotti et al. (2022). Faraday waves are striking standing wave patterns that emerge on a fluid surface subjected to parametric driving, i.e., when an external control parameter (like vertical oscillation amplitude) is varied periodically. First documented by Michael Faraday in 1831 Faraday (1831), this phenomenon represents a spontaneous breaking of both temporal and spatial symmetry. Significantly, Faraday waves have been investigated in diverse quantum systems Staliunas et al. (2002, 2004); Nicolin et al. (2007); Bhattacherjee (2008); Nath and Santos (2010); Nicolin and Raportaru (2010); Pollack et al. (2010); Nicolin (2011); Capuzzi et al. (2011); Balaž and Nicolin (2012); Łakomy et al. (2012); Raportaru (2012); Balaž et al. (2014); Abdullaev et al. (2015); Sudharsan et al. (2016); Nguyen et al. (2019); Vudragović and Balaž (2019); Abdullaev et al. (2019, 2019); Turmanov et al. (2020); Liu and Tu (2020); Cheng and Shi (2021); Zhang et al. (2020); Maity et al. (2020); Kwon et al. (2021); Zhang et al. (2022); Shukuno et al. (2023); Brito et al. (2023); Liang et al. (2024b); Wan et al. (2024); del R´ıo-Lima et al. (2024); Wang et al. (2025), including their first experimental demonstration in atomic BECs by Engels et al. Engels et al. (2007) and subsequent studies in dipolar condensates Łakomy et al. (2012); Turmanov et al. (2020), two-component BECs Chen et al. (2019); Wang et al. (2025), spin-orbit-coupled BECs Zhang et al. (2022) and Rydberg-dressed BECs Shi et al. (2023). These theoretical analyses and experiments establish that Faraday waves directly probe the underlying collective excitation spectrum of the quantum fluid. Therefore, driving Faraday waves in SGBECs offers a novel and controlled framework for investigating the collective dynamics, particularly the excitation spectrum of quantum fluids dominated by long-range attractive interactions.
In this paper, we present a theoretical investigation of Faraday waves in SGBECs, driven by periodic modulation of the $s$-wave scattering length. We employ linear stability analysis of the GPN equations to derive a Mathieu-like equation governing the parametric instability. This allows us to map out the stability phase diagram, revealing the interplay between the unstable tongues characteristic of parametric resonance and the Jeans instability inherent to self-gravitating systems. We analyze the dependence of the Faraday wavevector on the driving parameters and self-gravity. Furthermore, we incorporate dissipation effects and present numerical simulations illustrating the formation and evolution of Faraday waves within the SGBEC.
The paper is structured as follows: Section II outlines the theoretical model, the GPN equation with time-dependent interaction strength, and details the linear stability analysis leading to the Mathieu-like equation. Section III presents the resulting stability phase diagrams and analyzes the characteristics of the Faraday wavevector. Section IV discusses numerical simulations of Faraday wave dynamics within the driven SGBEC, including the effects of dissipation. We conclude in Section V.
II Model and Linear stability analysis
II.1 The Gross-Pitaevskii-Newton equation
The dynamics of self-gravitating Bose-Einstein condensates (SGBECs) is governed by the coupled Gross-Pitaevskii-Newton (Poisson) (GPN or GPP) equations:
$\mathrm{i}\hbar\frac{\partial\Psi}{\partial t}$ $=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\mathrm{ext}}\Psi+g(t)|\Psi|^{2}\Psi+m\Phi\Psi,$ (1a)
$\nabla^{2}\Phi$ $=4\pi Gm|\Psi|^{2}.$ (1b)
where $\Psi(\mathbf{r},t)$ is the condensate wave function, $V_{\rm ext}$ denotes the external potential, $m$ is the particle mass, and $\Phi(\mathbf{r},t)$ is the gravitational potential which satisfies the Poisson equation. The nonlinear coupling strength $g(t)=4\pi\hbar^{2}a_{s}(t)/m$ depends on the $s$-wave scattering length $a_{s}(t)$, tunable via Feshbach resonance. The GPN system can be expressed equivalently in integral form:
${\rm i}\hbar\frac{\partial\Psi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\rm ext}\Psi+g(t)|\Psi|^{2}\Psi-m^{2}G\int d\mathbf{r}^{\prime}\frac{|\Psi(\mathbf{r}^{\prime},t)|^{2}}{|\mathbf{r}-\mathbf{r}^{\prime}|}\Psi.$ (2)
The apparent singularity at $\mathbf{r}\to\mathbf{r}^{\prime}$ in the integral formulation is regularized by the finite extent of the wavefunction $|\Psi(\mathbf{r}^{\prime},t)|^{2}$, which acts as a natural cutoff. In practice, for a smooth, square-integrable wavefunction, this integral remains finite. The equivalence between the differential and integral forms can be rigorously established using distribution theory, where the Poisson equation $\nabla^{2}\Phi=4\pi Gm|\Psi|^{2}$ is interpreted in the weak sense, and the solution is given by the convolution with the Newtonian potential ${1}/{|\mathbf{r}-\mathbf{r}^{\prime}|}$, which is a fundamental solution of the Laplacian. Without gravity ($\Phi=0$), Eq. (1) reduces to the standard Gross-Pitaevskii equation; neglecting the mean-field term ($g=0$) yields the Schrödinger-Newton equation.
A generalized GPN model incorporating logarithmic nonlinearity and dissipation is proposed in Chavanis (2017):
${\rm i}\hbar\frac{\partial\Psi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\Psi+V_{\rm ext}\Psi+g(t)|\Psi|^{2}\Psi+m\Phi\Psi-\beta^{-1}\ln|\Psi|\Psi-{\rm i}\hbar\gamma\Psi,$ (3)
where $\gamma$ is the damping coefficient. The term $\beta^{-1}\ln|\Psi|\Psi$ accounts for finite-temperature effects, where $\beta^{-1}=-k_{B}T$ ($k_{B}$ is Boltzmann’s constant, $T$ is temperature), and the last term describes phenomenological dissipation. Eq. (1) is recovered when $\beta^{-1}=0$ and $\gamma=0$.
II.2 Reduction to Quasi-One-Dimensional System
We consider a cigar-shaped condensate confined by a fully anisotropic harmonic potential $V_{\rm ext}(\mathbf{r})=\frac{1}{2}m[\omega_{\perp}^{2}(x^{2}+y^{2})+\omega_{z}^{2}z^{2}]$, where $\omega_{\perp}$ and $\omega_{z}$ represent the transverse and longitudinal confinement frequencies, respectively. To ensure the system maintains a quasi-one-dimensional character, we require strong transverse confinement relative to the longitudinal direction, i.e., $\omega_{\perp}\gg\omega_{z}$. Under this condition, the transverse degrees of freedom are energetically frozen into the ground state, while the longitudinal motion remains dynamic. This allows for the factorization of the wave function as:
$\Psi(\mathbf{r},t)=\psi_{B}(\mathbf{r}_{\perp})\psi_{P}(z,t),$ (4)
where $\psi_{B}(\mathbf{r}_{\perp})$ is the transverse ground state, and $\psi_{P}(z,t)$ describes the longitudinal dynamics of the condensate. The transverse ground state is well approximated by a Gaussian:
$\psi_{B}(\mathbf{r}_{\perp})=\frac{1}{\sqrt{\pi}a_{\perp}}e^{-(x^{2}+y^{2})/(2a_{\perp}^{2})},\quad a_{\perp}=\sqrt{\frac{\hbar}{m\omega_{\perp}}}.$ (5)
Here, $a_{\perp}$ is the transverse harmonic oscillator length, which characterizes the size of the condensate in the transverse direction.
Substituting Eq.(4) into Eq.(3) and integrating over transverse coordinates, we obtain the effective quasi-1D equation(see Appendix A for details):
${\rm i}\hbar\frac{\partial\psi_{P}}{\partial t}=-\frac{\hbar^{2}}{2m}\frac{\partial^{2}\psi_{P}}{\partial z^{2}}+g_{\rm 1D}(t)|\psi_{P}|^{2}\psi_{P}+m\Phi_{\rm 1D}\psi_{P}-\beta^{-1}\ln|\psi_{P}|\psi_{P}-{\rm i}\hbar\gamma\psi_{P},$ (6)
where the effective 1D coupling is $g_{\rm 1D}(t)=g(t)/(2\pi a_{\perp}^{2})$ and $\Phi_{\rm 1D}$ is the quasi-1D gravitational potential. For small perturbations around the homogeneous background $\psi_{P}^{(0)}=\sqrt{n_{z}}$, where $n_{z}$ is the linear density (number of atoms per unit length), the self-gravity term can be approximated in the long-wavelength limit ($ka_{\perp}\ll 1$):
$m\Phi_{\rm 1D}\psi_{P}=-\frac{2Gm^{2}}{a_{\perp}}|\psi_{P}|^{2}\psi_{P}.$ (7)
This approximation is valid when the characteristic wavelength of density perturbations is much larger than the transverse confinement length scale $a_{\perp}$.
II.3 Linear stability analysis
Faraday waves in Bose-Einstein condensates can be excited through two primary methods: periodic modulation of the external trapping potential or time-dependent variation of the interaction strength via Feshbach resonance Engels et al. (2007). This work employs the latter approach, modulating the scattering length to produce a time-dependent quasi-1D coupling strength:
$g_{\rm 1D}(t)=g_{\rm 1D,0}\left(1+\alpha\cos\omega t\right),$ (8)
where $g_{\rm 1D,0}=g_{0}/(2\pi a_{\perp}^{2})=2\hbar^{2}a_{s}/(ma_{\perp}^{2})$ is the unmodulated quasi-1D coupling strength, $\omega$ is the driving frequency, and $\alpha$ is the modulation depth. We consider a cigar-shaped condensate tightly confined transversely, with driving applied along the longitudinal $z$-direction.We consider a cigar-shaped condensate tightly confined transversely, with driving applied along the longitudinal $z$-direction.
To analyze the stability against density modulations, we consider a plane wave perturbation with a general wavenumber $k$. The characteristic wavelength of such a modulation is $\lambda=2\pi/k$. The periodic driving will resonantly amplify a specific mode whose wavenumber $k_{F}$ (and corresponding wavelength $\lambda_{F}=2\pi/k_{F}$) satisfies the parametric resonance condition. When this Faraday wavelength $\lambda_{F}$ satisfies $\lambda_{F}\ll R_{\rm TF}$ (where $R_{\rm TF}$ is the Thomas-Fermi radius in the longitudinal direction), the external potential $V_{\rm ext}$ can be neglected, and the system can be treated as effectively homogeneous for the purpose of analyzing the Faraday wave formation.
The unperturbed state is homogeneous: $\psi_{P}^{(0)}=\sqrt{n_{z}}$, We introduce a perturbation:
$\psi_{P}(z,t)=\sqrt{n_{z}}\left[1+W(t)\cos(kz)\right],\quad W(t)=U(t)+{\rm i}V(t).$ (9)
[…]
$\hbar\dot{V}$ $=-\epsilon_{k}U-2g_{\mathrm{1D}}(t)\,n_{z}U+\frac{2m\Omega_{J,\mathrm{1D}}^{2}}{k^{2}}U+\beta^{-1}U-\hbar\gamma V,$ (10a)
$\hbar\dot{U}$ $=\epsilon_{k}V-\beta^{-1}V+\hbar\gamma U,$ (10b)
where $\epsilon_{k}=\hbar^{2}k^{2}/(2m)$ is the free-particle energy, and $\Omega_{J,{\rm 1D}}=\sqrt{4\pi Gm^{2}n_{2D}}$ is the quasi-1D Jeans frequency, with $n_{2D}=n_{z}/a_{\perp}$ being the effective areal density (number of atoms per unit area). The Jeans frequency derives its name from Sir James Jeans, who in 1902 first analyzed the gravitational instability of an infinite homogeneous medium in the context of astrophysics Jeans (1902). The Jeans criterion determines when self-gravity overcomes pressure support, leading to collapse Khlopov et al. (1985). In our quantum analog, $\Omega_{J,{\rm 1D}}$ represents the characteristic frequency at which self-gravity operates in the quasi-1D geometry, playing a role analogous to the classical Jeans frequency in astrophysical systems. For notational simplicity, we will hereafter denote $\Omega_{J,\mathrm{1D}}$ simply as $\Omega_{J}$ in the remainder of this paper.
Combining these equations gives a damped Mathieu equation:
$\frac{{\rm d}^{2}U}{{\rm d}t^{2}}+2\gamma\frac{{\rm d}U}{{\rm d}t}+\left[\Omega_{k}^{2}+\frac{\alpha g_{\rm 1D,0}n_{z}}{m}k^{2}\cos\omega t\right]U=0,$ (11)
with the characteristic excitation frequency
$\Omega_{k}^{2}=\Omega_{\rm Bog}^{2}-\Omega_{J}^{2}.$ (12)
Here $\Omega_{\rm Bog}$ is the Bogoliubov spectrum:
$\Omega_{\rm Bog}^{2}=\frac{\hbar^{2}k^{4}}{4m^{2}}+c_{s}^{2}k^{2},$ (13)
where the sound speed $c_{s}^{2}=\frac{g_{\rm 1D,0}n_{z}}{m}$. The derivation of Eq. (11) from the linearized Eqs. (10a) and (10b) employs two key approximations to arrive at the standard form of the damped Mathieu equation. First, the weak damping approximation ($\gamma\ll\omega,\Omega_{k}$) is applied, whereby terms of order $\gamma^{2}$ and higher are neglected. This is consistent with the phenomenological inclusion of dissipation and justifies the retention of only the first-order term $2\gamma\dot{U}$. Second, we consider the long-wavelength, low-temperature regime where the thermal energy scale satisfies $\beta^{-1}\ll\epsilon_{k}$. This ensures that the correction from the logarithmic nonlinearity remains small compared to the quantum pressure term, allowing us to cleanly isolate the parametric driving term. Under these physically reasonable conditions, the dynamics of the perturbation amplitude $U$ is faithfully described by the Mathieu equation, which captures the interplay between parametric resonance and Jeans instability.
Figure 1 illustrates the collective excitation spectrum of SGBECs. At large $k$ ($k\xi\gg 1$, where $\xi=\hbar/\sqrt{2mg_{\rm 1D,0}n_{z}}$ is the healing length), $\Omega_{k}$ approaches the Bogoliubov dispersion. The critical wavenumber $k_{J}$, indicated by the vertical dashed line in Fig. 1, marks the onset of Jeans instability and is defined by the condition $\Omega_{k}^{2}=0$. For $k<k_{J}$ (specifically, $k_{J}\xi\equiv\sqrt{\sqrt{1+\Omega^{2}_{J}/\Omega^{2}_{0}}-1}$, where $\Omega_{0}=g_{\rm 1D,0}n_{z}/\hbar$), the Jeans instability causes $\Omega_{k}^{2}<0$, indicating exponential growth. This critical wavenumber $k_{J}$ defines the quantum Jeans scale Chavanis (2020).
Figure 1: (Color online) Collective excitation spectrum of SGBECs. Parameters: $\Omega_{0}=g_{\rm 1D,0}n_{z}/\hbar$, $\xi=\hbar/\sqrt{2mg_{\rm 1D,0}n_{z}}$. The vertical dashed line indicates the critical wavenumber $k_{J}$ where $\Omega_{k}^{2}=0$. The region ($k<k_{J}$) indicates Jeans instability with $\Omega_{k}^{2}<0$.
Eq. (11) is rendered dimensionless via $\tau=\omega t/2$:
$\frac{{\rm d}^{2}U}{{\rm d}\tau^{2}}+\bar{\gamma}\frac{{\rm d}U}{{\rm d}\tau}+\left[\lambda-2q\cos(2\tau)\right]U=0,$ (14)
with parameters
$\lambda=\frac{4\Omega_{k}^{2}}{\omega^{2}},\quad q=-\frac{2\alpha g_{\rm 1D,0}n_{z}}{m}\frac{k^{2}}{\omega^{2}},$ (15)
and $\bar{\gamma}\equiv 4\gamma/\omega$.
In the dissipationless limit ($\gamma=0$), Eq. (14) is reduced to the canonical Mathieu equation, whose stability properties are well-established McLachlan (1947); Abramowitz and Stegun (1965); Wang and Guo (1989). The parameter $\lambda$ encodes the underlying excitation spectrum of the self-gravitating quantum fluid. We analyze its phase diagram and derive the Faraday wavevector in the following section.
III Stability and parametric resonance
III.1 Stability phase diagram
According to Floquet theory, the solution of Eq. (14) in the dissipationless limit can be expressed as:
$U(\tau)=e^{\mu\tau}f(\tau),$ (16)
[…]
Figure 2: (Color online) Stability phase diagram showing primary ($\nu=1$), secondary ($\nu=2$), and third ($\nu=3$) instability tongues. Shaded regions denote unstable solutions.
The numerically computed stability phase diagram is presented in Fig. 2. For small values of $q$, unstable regions emerge near $\lambda=\nu^{2}$ ($\nu=1,2,3,\ldots$), consistent with the characteristic instability tongues of parametric resonance. The primary instability tongue ($\nu=1$) exhibits the largest spatial extent in parameter space, with higher-order tongues ($\nu=2,3,\ldots$) progressively narrowing. For the dominant $\nu=1$ tongue ($\lambda=1$), the resonance condition corresponds to:
$\Omega_{k}=\omega/2,$ (17)
which yields the frequency-matching relation:
$\left(\frac{\omega}{2}\right)^{2}+\Omega_{J}^{2}=\Omega_{\rm Bog}^{2}.$ (18)
Eq.(18) indicates that the Faraday wave only occurs in the SGBEC when the relation Eq.(18) is met.
Figure 3 illustrates the stability phase diagrams at different Jeans frequencies. Without self-gravity ($\Omega_{J}=0$, Fig. 3(a)), instability arises exclusively from parametric driving. As $\Omega_{J}$ increases (Figs. 3(b)-(d)), the unstable region near $\lambda\approx 0$ expands significantly due to the relationship $\lambda=4\Omega_{k}^{2}/\omega^{2}$. When $\Omega_{\rm Bog}^{2}<\Omega_{J}^{2}$, $\lambda$ becomes negative and its magnitude scales as $|\lambda|\propto\Omega_{J}^{2}$, explaining the expansion of the Jeans-dominated region. Concurrently, the parametric resonance tongues shift toward larger $\lambda$ values. This shift occurs because $\lambda\propto k^{2}$ at intermediate wavevectors (where $\Omega_{k}^{2}\propto k^{2}$), and increased gravitational attraction requires shorter wavelengths to satisfy the parametric resonance condition $\lambda=\nu^{2}$, with the shift magnitude given by $\Delta\lambda\approx(4/\omega^{2})\Delta(\Omega_{J}^{2})$.
Figure 3: (Color online) Stability phase diagrams under different Jeans frequencies with $\Omega_{0}/\omega=0.5$. The region of Jeans instability ($\lambda<0$) is denoted by the label ”JI” in the figure.
Figure 4 illustrates the central result of our stability analysis: the driving frequency $\omega$ serves as a key control parameter governing the competition between intrinsic Jeans instability and driven parametric resonance.
Figure 4: (Color online) Stability phase diagrams for a fixed Jeans frequency $\Omega_{J}/\Omega_{0}=1$ at progressively lower driving frequencies $\omega$, showing the expansion of the Jeans-unstable region ($\lambda<0$, labeled as ”JI”) and the concomitant fragmentation and narrowing of the parametric resonance tongues. The clear transition from a parametric-resonance-dominated regime to a Jeans-dominated regime is evident.
At low driving frequencies ($\omega\ll\Omega_{0},\Omega_{J}$), the stability diagram is dominated by gravity-induced instability. The Jeans-unstable region (labeled ”JI” in Fig. 4) expands significantly, as $|\lambda|\propto\Omega_{J}^{2}/\omega^{2}$ becomes large. Consequently, the tongues associated with parametric resonance become increasingly fragmented, narrowed, and eventually overwhelmed by the expansive domain of the intrinsic Jeans instability. In this regime, the system’s dynamics are governed by gravitational collapse, and the formation of distinct Faraday wave patterns is suppressed.
Conversely, at high driving frequencies ($\omega\gg\Omega_{0},\Omega_{J}$), the influence of the Jeans instability is markedly reduced. The Jeans-unstable region (labeled ”JI”) is compressed into a narrow band near $\lambda\approx 0$ due to the scaling $\lambda\propto 1/\omega^{2}$. Simultaneously, the parametric resonance tongues shift to larger values of $\lambda$ and, crucially, become well-separated from each other and from the Jeans band. In this regime, the instability landscape is dominated by these distinct parametric resonance structures, which define clear and accessible parameter windows for the formation of Faraday waves.
The most dramatic interplay between these mechanisms occurs for driving frequencies near the characteristic system frequency ($\omega\sim\Omega_{0}$). Here, the parametric resonance tongues undergo significant fragmentation and narrowing, governed by the scalings $\Delta\lambda\sim\omega^{-2}$ and $\Delta q\sim\omega^{-4}$. This complex transition region showcases the direct competition between the two instability mechanisms.
Thus, increasing the driving frequency does not eliminate the fundamental Jeans instability but effectively suppresses its relative dominance in the parameter space. This suppression creates a well-defined regime where the driven, parametric instability becomes the dominant and observable outcome, enabling the controlled generation of Faraday waves. This tunability-from gravity-dominated collapse (as seen in the expanding ”JI” regions at low frequencies) to pattern-forming resonance-underscores the utility of periodic driving for controlling nonlinear instabilities in self-gravitating quantum fluids and reveals a rich phase diagram structure that could be exploited for precise dynamical control.
III.2 The Faraday wavevector
From the instability condition $\lambda=\nu^{2}$ for the $\nu^{\text{th}}$ resonance tongue, we obtain the Faraday wavevector $k_{F}$:
$k_{F}\xi=\sqrt{\sqrt{1+\frac{1}{\Omega_{0}^{2}}\left(\Omega_{J}^{2}+\frac{\nu^{2}\omega^{2}}{4}\right)}-1}.$ (19)
In the absence of self-gravity ($\Omega_{J}=0$), this expression reduces to the well-known result for conventional BECs Hernández-Rajkov et al. (2021):
$k_{F}^{(0)}\xi=\sqrt{\sqrt{1+\left(\frac{\nu\omega}{2\Omega_{0}}\right)^{2}}-1}.$ (20)
Figure 5 shows $k_{F}$ as a function of driving frequency $\omega$ for the first three branches. The Faraday wavevector increases with driving frequency, approaching the behavior of non-self-gravitating BECs at large $\omega$. At low driving frequencies, $k_{F}\xi$ approaches a finite value $\sqrt{\sqrt{\Omega_{J}^{2}/\Omega_{0}^{2}+1}-1}$ in SGBECs, contrasting with the $k_{F}\rightarrow 0$ behavior in non-gravitational systems. Higher $\nu$ values yield larger $k_{F}$ at fixed $\omega$, consistent with excitation of shorter-wavelength modes.
Figure 5: (Color online) Faraday wavevector versus driving frequency for $\nu=1,2,3$ with $\Omega_{J}/\Omega_{0}=1.2$.
IV Numerical results
Our numerical simulations are based on solving the dimensionless damped Mathieu equation governing the perturbation amplitude $U(\tau)$ with fixed parameters $\lambda=1$ and q=0.1q=-0.1, where $\bar{\gamma}\equiv 4\gamma/\omega$ is the dimensionless damping coefficient. This configuration satisfies the primary parametric resonance condition $2\Omega_{k}=\omega$ for $\nu=1$. Fig. 6 presents the dynamical evolution across four dissipation strengths, with each column corresponding to: (a) $\bar{\gamma}=0$ exhibiting exponential growth; (b) $\bar{\gamma}=0.01$ showing persistent instability; (c) $\bar{\gamma}=0.02$ demonstrating suppressed growth; and (d) $\bar{\gamma}=0.05$ displaying overdamped decay. The top row tracks $U(\tau)$, while the middle row shows background-subtracted density $|\Psi_{P}|^{2}/n_{z}-1$ to isolate Faraday waves from the uniform ground state, and the bottom row displays momentum-space density.
Figure 6: (Color online) Density wave dynamics under dissipation strengths $\bar{\gamma}=0$ (a), $0.01$ (b), $0.02$ (c), $0.05$ (d) with $\lambda=1$, q=0.1q=-0.1.
For $\bar{\gamma}=0$ (Fig. 6(a)), parametric instability drives exponential growth of $U(\tau)$ with Floquet exponent $\mu\approx 0.2$, generating pronounced density waves of wavelength $\lambda_{F}=2\pi/k$ and sharp momentum-space peaks at $\pm k$. At $\bar{\gamma}=0.01$ (Fig. 6(b)), dissipation fails to suppress the instability. $U(\tau)$ maintains strong exponential growth with minimal amplitude reduction. This confirms that weak dissipation cannot overcome parametric resonance. At $\bar{\gamma}=0.02$ (Fig. 6(c)), the competition between dissipation and parametric driving becomes evident: while unstable evolution persists, amplitude growth is significantly suppressed, resulting in weak density modulations. Under strong dissipation ($\bar{\gamma}=0.05$, Fig. 6(d)), dissipation dominates, causing rapid decay of $U(\tau)$ within $\tau<20$ and complete suppression of density waves. This sequence illustrates the transition from parametric-resonance-dominated ($\bar{\gamma}<0.02$) to dissipation-dominated ($\bar{\gamma}>0.02$) regimes. Gravitational effects influence this competition through the $\lambda$-parameter in Eq. (14), which encodes the $\Omega_{J}$-modified excitation spectrum.
In Figure 7, we present the evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies $\Omega_{J}/\Omega_{0}=0,0.1,0.5,1.2$, with fixed parameters $\omega/\Omega_{0}=1$, q=0.1q=-0.1, and $\bar{\gamma}=0.01$. The corresponding $\lambda$ values, determined by the relation $\lambda=1-4(\Omega_{J}/\Omega_{0})^{2}$, are $\lambda=1.00$, $0.96$, $0.00$, and 4.76-4.76 respectively. For no self-gravity ($\Omega_{J}/\Omega_{0}=0$, $\lambda=1$), the system resides at the center of the primary parametric resonance tongue, resulting in exponential growth of $U(\tau)$ despite weak damping ($\bar{\gamma}=0.01$), characteristic of parametric instability that seeds Faraday waves. As $\Omega_{J}/\Omega_{0}$ increases to $0.1$ ($\lambda=0.96$), the system remains within the parametric resonance tongue but approaches its boundary, exhibiting sustained growth of $U(\tau)$ at a reduced rate compared to the $\Omega_{J}=0$ case, with Faraday waves forming at diminished amplitudes due to competition between parametric driving and incipient gravitational effects. At the critical transition point ($\Omega_{J}/\Omega_{0}=0.5$, $\lambda=0$) where $\Omega_{k}=0$, parametric resonance and Jeans instability coexist, leading to irregular, non-exponential growth of $U(\tau)$ that deviates from characteristic Faraday patterns. Finally, in the Jeans-dominated regime ($\Omega_{J}/\Omega_{0}=1.2$, $\lambda=-4.76$), self-gravity overwhelms parametric driving ($\Omega_{k}^{2}<0$), causing rapid exponential growth of $U(\tau)$ characteristic of gravitational collapse that suppresses Faraday wave formation. This progression confirms the phase diagram prediction (Fig. 4) where increasing Jeans instability fragments parametric resonance structures, with the transition at $\Omega_{J}/\Omega_{0}\sim 0.5$ ($\lambda=0$) highlighting the sensitivity of driven quantum fluids to long-range attractive interactions.
Figure 7: Evolution of the perturbation amplitude $U(\tau)$ under varying Jeans frequencies: (a) $\Omega_{J}/\Omega_{0}=0$ ($\lambda=1$), (b) $\Omega_{J}/\Omega_{0}=0.1$ ($\lambda=0.96$), (c) $\Omega_{J}/\Omega_{0}=0.5$ ($\lambda=0$), (d) $\Omega_{J}/\Omega_{0}=1.2$ ($\lambda=-4.76$). Fixed parameters: $\omega/\Omega_{0}=1$, q=0.1q=-0.1, $\bar{\gamma}=0.01$. Exponential growth in (a) signifies parametric instability, while rapid growth in (d) reflects Jeans collapse.
The numerical evolution of the perturbation amplitude $U(\tau)$ in Fig. 7 provides a temporal signature of the competition and transition between instability mechanisms. The progression from clear oscillatory growth (parametrically driven Faraday waves) at $\lambda=1$ to pure monotonic growth (Jeans collapse) at $\lambda=-4.76$, with a crossover region of irregular growth at $\lambda=0$, corroborates the phase diagram prediction. It visually demonstrates that driving competes with the intrinsic instability, and the final outcome is determined by their relative strength in the $\lambda$-$q$ parameter space.
V Conclusion
In this work, we have demonstrated that the stability landscape of a driven SGBEC is characterized by the competitive interplay between its intrinsic Jeans instability and driven parametric resonance. The central finding is that this competition can be effectively controlled and tuned via the driving frequency. High-frequency driving suppresses the prominence of Jeans instability in parameter space and enables the domination of parametric resonance, creating a well-defined regime for probing collective excitations via Faraday waves. Our work thus establishes periodic driving not as a way to eliminate fundamental instabilities, but as a tool to manipulate their competition and access novel non-equilibrium states in quantum fluids with long-range interactions.
Notably, while this study employed modulation of the scattering length, Faraday waves may also be excited through periodic modulation of the external potential in similar configurations. Given the significance of the generalized GPN equations in astrophysical contexts, our approach provides a flexible platform for exploring gravitational analogs. Specifically, such systems could model scenarios where external potential fields originate from compact objects like black holes, and coupling constants emerge from self-gravitating dark matter configurations. This framework offers promising pathways for laboratory investigations into gravitational phenomena using quantum degenerate gases.
Acknowledgments
This work was supported by the Open Fund of Key Laboratory of Multiscale Spin Physics (Ministry of Education), Beijing Normal University (Grant No.SPIN2024N03), and the Scientific Research Startup Foundation for High-Level Talents at Anqing Normal University (Grant No.241042). We thank Professor Zhanchun Tu for valuable discussions and insightful comments.
Appendix A Reduction to quasi-1D equations for self-gravitational potential
The 3D gravitational potential is given by the solution of the Poisson equation:
$\Phi(\mathbf{r},t)=-Gm\int\frac{|\Psi(\mathbf{r}^{\prime},t)|^{2}}{|\mathbf{r}-\mathbf{r}^{\prime}|}d^{3}\mathbf{r}^{\prime}.$ (21)
Substituting the factorized wave function:
$\Phi(\mathbf{r},t)=-Gm\int dz^{\prime}|\psi_{P}(z^{\prime},t)|^{2}\int d^{2}\mathbf{r}_{\perp}^{\prime}\frac{|\psi_{B}(\mathbf{r}_{\perp}^{\prime})|^{2}}{|\mathbf{r}-\mathbf{r}^{\prime}|}.$ (22)
The quasi-1D potential is defined as the transverse average:
$\Phi_{\rm 1D}(z,t)=\int|\psi_{B}(\mathbf{r}_{\perp})|^{2}\Phi(\mathbf{r},t)d^{2}\mathbf{r}_{\perp}.$ (23)
Substituting $\Phi(\mathbf{r},t)$:
$\Phi_{\rm 1D}(z,t)=-Gm\int dz^{\prime}K(z-z^{\prime})|\psi_{P}(z^{\prime},t)|^{2}.$ (24)
The integral over $\mathbf{r}_{\perp}$ and $\mathbf{r}_{\perp}^{\prime}$ defines the kernel:
$K(z-z^{\prime})=\int d^{2}\mathbf{r}_{\perp}|\psi_{B}(\mathbf{r}_{\perp})|^{2}\int d^{2}\mathbf{r}_{\perp}^{\prime}\frac{|\psi_{B}(\mathbf{r}_{\perp}^{\prime})|^{2}}{\sqrt{|\mathbf{r}_{\perp}-\mathbf{r}_{\perp}^{\prime}|^{2}+(z-z^{\prime})^{2}}}.$ (25)
Using the Fourier representation of the potential:
$\frac{1}{|\mathbf{r}|}=\frac{1}{2\pi^{2}}\int\frac{e^{i\mathbf{k}\cdot\mathbf{r}}}{k^{2}}d^{3}k,$ (26)
we can write:
$\frac{1}{\sqrt{|\mathbf{r}_{\perp}-\mathbf{r}_{\perp}^{\prime}|^{2}+(z-z^{\prime})^{2}}}=\frac{1}{2\pi^{2}}\int\frac{e^{i\mathbf{k}_{\perp}\cdot(\mathbf{r}_{\perp}-\mathbf{r}_{\perp}^{\prime})+ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}d^{2}k_{\perp}dk_{z}.$ (27)
Substituting this into $K(z-z^{\prime})$ and using the Gaussian form of $|\psi_{B}|^{2}$, we obtain:
$K(z-z^{\prime})=\frac{1}{2\pi^{2}}\int d^{2}k_{\perp}dk_{z}\frac{e^{ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}\left|\int d^{2}\mathbf{r}_{\perp}|\psi_{B}(\mathbf{r}_{\perp})|^{2}e^{i\mathbf{k}_{\perp}\cdot\mathbf{r}_{\perp}}\right|^{2}.$ (28)
The Gaussian integral evaluates to:
$\int d^{2}\mathbf{r}_{\perp}|\psi_{B}(\mathbf{r}_{\perp})|^{2}e^{i\mathbf{k}_{\perp}\cdot\mathbf{r}_{\perp}}=e^{-k_{\perp}^{2}a_{\perp}^{2}/4},$ (29)
so:
$K(z-z^{\prime})=\frac{1}{2\pi^{2}}\int d^{2}k_{\perp}dk_{z}\frac{e^{ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}e^{-k_{\perp}^{2}a_{\perp}^{2}/2}.$ (30)
Performing the $k_{z}$ integration first via contour integration:
$\int_{-\infty}^{\infty}\frac{e^{ik_{z}(z-z^{\prime})}}{k_{\perp}^{2}+k_{z}^{2}}dk_{z}=\frac{\pi}{k_{\perp}}e^{-k_{\perp}|z-z^{\prime}|},$ (31)
we get:
$K(z-z^{\prime})=\frac{1}{2\pi}\int_{0}^{\infty}k_{\perp}dk_{\perp}\frac{e^{-k_{\perp}|z-z^{\prime}|}}{k_{\perp}}e^{-k_{\perp}^{2}a_{\perp}^{2}/2}=\frac{1}{2\pi}\int_{0}^{\infty}e^{-k_{\perp}|z-z^{\prime}|}e^{-k_{\perp}^{2}a_{\perp}^{2}/2}dk_{\perp}.$ (32)
This integral can be evaluated exactly and yields:
$K(z-z^{\prime})=\frac{1}{\sqrt{2\pi}a_{\perp}}e^{(z-z^{\prime})^{2}/2a_{\perp}^{2}}\Gamma\left(\frac{1}{2},\frac{(z-z^{\prime})^{2}}{2a_{\perp}^{2}}\right),$ (33)
where $\Gamma(a,z)$ is the incomplete Gamma function. This expression is identically equal to:
$K(z-z^{\prime})=2\pi K_{0}\left(\frac{|z-z^{\prime}|}{\ell}\right),$ (34)
where $K_{0}$ is the modified Bessel function of the second kind and $\ell=a_{\perp}/\sqrt{2}$. Thus:
$\Phi_{\rm 1D}(z,t)=-2Gm\int_{-\infty}^{\infty}dz^{\prime}\,|\psi_{P}(z^{\prime},t)|^{2}\,K_{0}\left(\frac{|z-z^{\prime}|}{\ell}\right).$ (35)
In the long-wavelength limit ($k\ell\ll 1$), this reduces to a local form:
$\Phi_{\rm 1D}(z,t)\approx-\frac{2Gm}{a_{\perp}}|\psi_{P}(z,t)|^{2}.$ (36)
This local approximation is valid when the characteristic wavelength of density perturbations is much larger than the transverse confinement length scale $\ell$.